8/[ ] = 3.2/4 8/[ ] = 3.2/4 8/[ ] = 3.2/4 8/[ ] = 3.2/4

 8/[ ] = 3.2/4
 8/[ ] = 3.2/4
 8/[ ] = 3.2/4
 8/[ ] = 3.2/4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the missing number be y.
Then, 8/y = 3.2/4
By cross multiplication we get,
y = (8 × 4)/3.2
y = 32/3.2
y = 320/32
Divide both numerator and denominator by 32.
y = 10
Therefore, 8/[10] = 3.2/4

Thanks